OpenStudy (anonymous):

What is the value of m? 14+2m=4m-16

OpenStudy (anonymous):

subtract 2m from both sides and add 16 to both sides: 14+2m-2m=4m-2m-16 14=2m-16 14+16=2m-16+16 30=2m divide both sides by 2 2/2m=30/2 m=15

OpenStudy (mrnood):

you need to get all the terms with m in on one side of the equation, and all the other terms on the other side so start by subtracting 2m from each side of the equaition or just take the complete answer that someone else posted and don't learn anything

OpenStudy (anonymous):

thx bro

OpenStudy (anonymous):

np :)